OpenStudy (anonymous):

solve the eniquality. x^3-x greater than or equal to 0

OpenStudy (anonymous):

this is the first step: x^3 - x = x(x^2 - 1) = x(x + 1)(x - 1)

OpenStudy (anonymous):

x(x+1)(x-1) >= 0

OpenStudy (anonymous):

eigenschmeigen you can also help me in a min too mate

OpenStudy (anonymous):

now we have this product. first consider the roots: x = -1, x = 0 , x= 1 now consider the cases x=<-1 , -1=<x=<0 , 0=<x=<1 , 1=<x

OpenStudy (anonymous):

oh ok so once you distribute you just set it equal to zero and find them?

OpenStudy (anonymous):

yes. in fact, most of this is unnecessary, but its good practice for harder problems drawing a graph helps: